# Few-shot Tuning of Foundation Models for Class-incremental Learning

### v1

@@ Abstract
For the first time, we explore few-shot tuning of vision foundation models for class-incremental learning. Unlike existing few-shot class incremental learning (FSCIL) methods, which train an encoder on a base session to ensure forward compatibility for future continual learning, foundation models are generally trained on large unlabelled data without such considerations. This renders prior methods from traditional FSCIL incompatible for FSCIL with the foundation model. To this end, we propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a new approach to continually tune foundation models for new classes in few-shot settings. CoACT comprises three components: (i) asynchronous contrastive tuning, which learns new classes by including LoRA modules in the pre-trained encoder, while enforcing consistency between two asynchronous encoders; (ii) controlled fine-tuning, which facilitates effective tuning of a subset of the foundation model; and (iii) consistency-guided incremental tuning, which enforces additional regularization during later sessions to reduce forgetting of the learned classes. We perform an extensive study on 16 diverse datasets and demonstrate the effectiveness of CoACT, outperforming the best baseline method by 2.47% on average and with up to 12.52% on individual datasets. Additionally, CoACT shows reduced forgetting and robustness in low-shot experiments. As an added bonus, CoACT shows up to 13.5% improvement in standard FSCIL over the current SOTA on benchmark evaluations. We make our code publicly available at https://github.com/ShuvenduRoy/CoACT-FSCIL.
@@ 1 Introduction
Few-Shot Class-Incremental Learning (FSCIL) is a continual learning paradigm that involves sequentially learning a set of new classes with only a limited number of samples available per class [51; 44; 39]. This task is particularly challenging yet important as it mimics real-world scenarios where new classes may emerge over time, and collecting a large training set for these new classes can be difficult. In practice, FSCIL assumes the availability of a large base session where the number of classes and the number of samples per class are typically large [51]. The base session is subsequently followed by incremental learning sessions, which introduce new classes with a limited number of samples per class. In the traditional paradigm of FSCIL, the model is usually trained on the base session with fully supervised learning, while facilitating forward compatibility for learning future incremental classes [42; 21]. Techniques for forward compatibility during base training include allocating space in the embedding space [42] or identifying specific parameters of the encoder that it deems necessary to keep frozen while learning the new sessions [21].
Figure 1: Traditional FSCIL vs. FSCIL with foundation models.
With the rise of large foundation models with excellent generalization capabilities and impressive performances across different domains, the ability to continually learn new classes with very little data can even further enhance their adaptability and effectiveness in real-world applications. However, foundation models are generally trained on web-scale unlabeled data without any knowledge (or compatibility) for future continual learning. This in turn renders existing FSCIL solutions impractical for FSCIL with a foundation model and requires a fresh approach where we aim to tune off-the-shelf foundation models for new classes without losing previously learned knowledge (see Figure 1). As far as we are aware, enabling FSCIL by tuning large foundation models has not been explored in prior works.
Tuning foundation models also comes with its own challenges, particularly when dealing with limited labelled data. Despite advancements in parameter-efficient tuning [23; 13] and regularization techniques [38], fine-tuning of foundation models in a few-shot setting often leads to a decline in generalization due to overfitting, which can result in catastrophic forgetting in the context of continual learning.
To address this challenge, we propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a novel framework for class-incremental learning in foundation models in few-shot settings. CoACT comprises three components: (i) asynchronous contrastive tuning, (ii) controlled fine-tuning, and (iii) consistency-guided incremental tuning. Asynchronous contrastive tuning learns from the first incremental session using a novel asynchronous contrastive method that strikes a balance between adaptability to learn new classes and retaining generalizable knowledge of the pre-trained foundation model. For adapting the model to the first new session, we integrate learnable LoRA modules into the pre-trained encoder, and ensure generalization by enforcing consistency between two asynchronous encoders: a student encoder containing the learnable modules, and a teacher encoder identical to the pre-trained encoder, updated as the Exponential Moving Average (EMA) of the student. This prevents rapid change in the output distribution of the teacher, which in turn helps reduce overfitting in the student encoder while learning effective representations. In addition to training LoRA, to further enhance adaptability, we introduce controlled fine-tuning, which utilizes a two-step training protocol for the first incremental session. First, we train the newly added LoRA modules with a high Learning Rate (LR) for a certain number of epochs and then fine-tune the last few layers of the pre-trained parameters with a relatively lower LR. This also helps balance adaptability with generalizability. Finally, consistency-guided incremental tuning is a novel regularization technique that ensures effective learning of classes in the following incremental sessions while preventing forgetting of previously learned classes and preserving the generalization capabilities of the foundation model at the same time. To achieve this, we enforce consistency between the predictions of the learnable encoder in the incremental sessions and the frozen encoder from the first incremental session.
We conduct a comprehensive study on 16 diverse image recognition datasets to investigate the effectiveness of our method. The datasets include generic objects, fine-grained objects, scenes, satellite images, and texture recognition. Considering that traditional FSCIL methods are not compatible for FSCIL with foundation models, we first establish a set of baselines by building on the few-shot learning literature, namely prototype-learning [42], linear tuning, and LoRA tuning [19]. Our comprehensive experiments demonstrate that CoACT achieves a 2.47% average improvement over the best baseline method, with up to 12.79% performance gain on individual datasets (Figure 3). More importantly, CoACT exhibits reduced forgetting of already learned classes as the number of classes increases. We also evaluate the efficacy of CoACT in a traditional FSCIL setup, where the first incremental session is considerably large. Following established protocols, we report these results on CIFAR-100, CUB-200, and miniImageNet datasets. Our method achieves state-of-the-art performance (Table 3), outperforming existing approaches by 4.11%, 5.02%, and 13.47% on the three datasets, respectively. We provide detailed ablation studies showing the effectiveness of each component of our method. Overall, we make the following contributions:
  • For the first time, we explore FSCIL with foundation models to enable continuous tuning of the model to learn new classes in few-shot settings.
  • We propose CoACT, a novel method for that can effectively perform FSCIL with foundation models without losing its generalization capability or forgetting already learned classes. Our method consists of three key components that enrich our repertoire for tuning foundation models for more dynamic real-world settings.
  • Comprehensive experiments show the effectiveness of our method, achieving SOTA on both traditional FSCIL and FSCIL with foundation models. We also show reduced forgetting and effectiveness in very-low-shot settings. Extensive ablation and sensitivity studies show the effectiveness of each of our proposed components and their optimal setup. We make our code publicly available: https://github.com/ShuvenduRoy/CoACT-FSCIL.
@@
Figure 2: Performance comparison on FSCIL with a foundation model.
@@
CoACT (Ours) 84.62 81.19 96.24
Figure 3: Performance comparison on tradition FSCIL.
@@ 2 Related works
Few-shot class-incremental learning. Class-incremental learning is a continual learning process that focuses on continuous learning of new classes while retaining the knowledge of already learned ones [30]. In practice, machine learning models often need to learn new classes from a few labelled samples per class [51], while having no access to samples from already learned classes. This scenario has given rise to a new learning task called few-shot class-incremental learning or FSCIL [44]. The existing literature on FSCIL can be broadly categorized into two main groups: methods that continuously train both the encoder and classifier over each incremental session  [43; 6; 10; 49], and methods that keep the encoder frozen during the incremental learning sessions [52; 41; 48]. We provide a brief summary of both groups below.
As an example of the first group of FSCIL methods, MgSvF [49] employed a component-wise update strategy to ensure adapting to new classes while preserving knowledge of existing ones. The exemplar relation distillation framework [10] constructed and updated a graph of exemplar relationships to facilitate the integration of new classes. SoftNet [21] introduced a novel approach for identifying and freezing a sub-network of crucial parameters from the previous session and training the remaining parameters during incremental sessions. While such methods in the first group generally offer greater adaptability to new classes compared to the second group: (a) they require additional constraints to avoid overfitting to the new classes and thus catastrophic forgetting; (b) they are not generally feasible in the context of foundation models as encoders are quite large, resulting in strong overfitting in few-shot settings.
In the second group of FSCIL methods, FACT [50] introduced virtual prototypes to maximize class embedding separation while preserving their relative positions. SAVC [42] generated virtual classes during the base session training to maximize separability. Similarly, NC-FSCIL [47] pre-assigned optimally spaced prototypes to each base class, promoting diverse and distinct class representations. Methods in this group generally focus on ensuring forward compatibility by maintaining separability in the base classes in the learned embedding space to enable the effective integration of new classes without interfering with existing ones [52; 41; 48]. While such methods in the second group generally offer greater stability on already learned classes: (a) stability often comes at the cost of adaptability toward learning new classes; (b) these methods are generally not compatible with foundation models given lack of control over the base training of such off-the-shelf models.
@@
Fine-tuning foundation models. A number of techniques have recently been proposed to tune foundation models without the need to re-train them from scratch. Adapter tuning [18] fine-tunes large pre-trained models to downstream tasks by inserting new learnable layers inside the pre-trained model. On the other hand, prompt-tuning [27] and prefix-tuning [28] add learnable prompts with the input embedding for learning the new task without tuning the pre-trained parameters of the model. Low-rank adapters [19; 22] have been introduced to reduce the computation cost of additional parameters. VPT [20] and AdapterFormer [4] explored parameter-efficient fine-tuning in the context of vision transformers. Nonetheless, such methods are not designed for continuous tuning of a foundation model since there are no inherent mechanisms in them to prevent loss of generalization capabilities and catastrophic forgetting.
@@ 3.1 Problem formulation
In FSCIL, a model $\phi(x)$ is continuously trained over $T$ consecutive sessions of new classes. Each session follows an $n$-way, $k$-shot setup, where $n$ represents the number of classes in each session, and $k$ is the number of samples per class. Consequently, the training data for each session $t\in T$ can be represented as $\mathcal{D}_{train}^{t}=\{(x_{i},~{}y_{i})\}_{i=1}^{N_{t}}$, where $x_{i}$ and $y_{i}$ denote the $i^{\text{th}}$ sample and its corresponding label, and $N_{t}$ represents the number of samples in session $t$. We can express $\phi(x)=W^{T}f_{\theta}(x)$, where $f_{\theta}$ is an encoder and $W$ is a linear classifier. In the conventional setup of FSCIL, $f_{\theta}$ is generally trained on a large labelled base session ($\mathcal{D}_{train}^{0}$) while ensuring forward compatibility toward learning new classes, and remains frozen in the incremental sessions. If we assume $f_{\theta}$ to be a foundation model pre-trained on large web-scale data with no knowledge of future continual learning, standard FSCIL techniques are rendered incompatible since they require a labelled base session. In this work, we aim to design a method for tuning off-the-shelf pre-trained foundation models to continuously learn new classes in few-shot learning settings.
@@ 3.2 Consistency-guided Asynchronous Contrastive Tuning (CoACT)
Figure 4: Illustration of CoACT. (Left) Training on the first incremental session with asynchronous contrastive tuning and controlled fine-tuning. The student encoder contains learnable LoRA modules, while the teacher is identical to the foundation model but updated as the EMA of the student. Controlled fine-tuning enables the tuning of a subset of the foundation model with reduced LR after certain epochs. (Right) Consistency-guided incremental tuning enforces consistency between the learnable student and the frozen encoder from the first incremental session, providing additional regularization that prevents overfitting and forgetting.
3.2.1 Asynchronous contrastive tuning
To strike a balance between adaptability to new classes and retaining generalizable knowledge of the foundation model, we introduce asynchronous contrastive tuning as the first component in our framework. This involves fine-tuning the pre-trained model using our novel Asynchronous Contrastive Learning (ACL) approach while incorporating LoRA modules into the foumodel. Let $h_{i}=f_{\theta}^{(i)}(h_{i-1})$ be the output of $i^{\text{th}}$ layer of the pre-trained encoder, $h_{i-1}$ be the output of the $(i-1)^{\text{th}}$ hidden layer of the encoder, and $f_{\theta}^{(i)}$ be the $i^{\text{th}}$ layer of the encoder. With the new learnable LoRA layers, the output of the $i^{\text{th}}$ layer of the network can be represented as $h_{i}^{{}^{\prime}}=f_{\theta}^{(i)}(h_{i-1})+f^{(i)}_{\text{LoRA}}(h_{i-1})$, where, $f^{(i)}_{\text{LoRA}}$ is the $i^{\text{th}}$ LoRA layer added to the pre-trained encoder. For brevity, we denote the encoder with learnable LoRA layers as $f_{\theta^{{}^{\prime}}}$. We can train $f_{\theta^{{}^{\prime}}}$ on $\mathcal{D}_{train}^{1}$ to learn the first incremental session as:
@@ 3.2.1 Asynchronous contrastive tuning
However, it has been shown in prior work that cross-entropy alone does not learn a well-separable embedding space [42] and has a higher over-fitting tendency, especially in a few-shot setting [38]. To reduce the possibility of overfitting and retain the generalization in the learnable encoder $f_{\theta^{{}^{\prime}}}$, we regularize its output distribution with a teacher encoder by maximizing their agreement in the embedding space. Here, the teacher encoder is identical to the pre-trained model (does not contain the LoRA modules) and learned through the Exponential Moving Average (EMA) of the student $f_{\theta^{{}^{\prime}}}$ as $\theta^{{}^{\prime\prime}}=m\cdot\theta^{{}^{\prime\prime}}+(1-m)\cdot\theta$, where $\theta$ and $\theta^{{}^{\prime\prime}}$ are the parameters of the student (without LoRA) and teacher encoders, and $m$ is the momentum parameter. Given that the teacher and student encoders differ in their architecture due to the addition of LoRA to the student, learning occurs asynchronously. This asynchronous encoder design and slow-moving update of the teacher through EMA ensures that the predictions from the teacher do not fluctuate. Since the teacher encoder is also initialized from the foundation model, consistency with the teacher effectively regularizes the student from overfitting.
In practice, we maximize the agreement between the embeddings of the student and teacher encoders on all the samples from each class using a supervised contrastive loss [24] as:
@@
3.2.2 Controlled fine-tuning
@@
To further enhance the adaptability of the model, we enable controlled fine-tuning of some of the pre-trained layers of the encoder. Since the newly added LoRA modules are randomly initialized, we begin by training only the LoRA modules of the student encoder with a higher learning rate for the initial $E_{c}$ epochs of training. This is followed by a fine-tuning stage where the last $C_{l}$ layers of the pre-trained encoder are fine-tuned with a reduced LR (scaled by a factor of $C_{f}$). We focus on fine-tuning only the last $C_{l}$ layers, as the later layers of a pre-trained model are responsible for learning domain-specific fine-grained features, whereas the earlier layers are more general and transferable to a wide range of tasks [32].
3.2.3 Consistency-guided incremental tuning
@@
While the first two modules facilitate tuning the foundation model ($f_{\theta^{{}^{\prime}}}$) during the first session, later sessions also require the retention of previously learned classes. To facilitate this, we propose consistency-guided incremental tuning, which prevents forgetting by regularizing the output distribution of the student $f_{\theta^{{}^{\prime}}}$ when training on the incremental sessions. More specifically, we enforce consistency between the predictions of the student encoder and a frozen encoder after the first session, effectively discouraging substantial changes in the learned representations of the student. Let $f_{\theta_{\beta}}$ be the frozen encoder after the first session, and the frozen embedding of this encoder be $p_{i}=f_{\theta_{\beta}}(\mathcal{A}(x_{i}))$. We define our consistency regularizer as:

### v2

# Consistency-Guided Asynchronous Contrastive Tuning for Few-Shot Class-Incremental Tuning of Foundation Models

@@ Abstract
We propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a novel method for continuously tuning foundation models to learn new classes in few-shot settings. CoACT consists of three key components:(i) asynchronous contrastive tuning, which learns new classes by including LoRA modules in the pre-trained encoder while enforcing consistency between two asynchronous encoders; (ii) controlled fine-tuning, which facilitates effective tuning of a subset of the foundation model; and (iii) consistency-guided incremental tuning, which enforces additional regularization during later sessions to reduce forgetting of the learned classes. We evaluate our proposed solution on Few-Shot Class-Incremental Learning (FSCIL) as well as a new and more challenging setup called Few-Shot Class-Incremental Tuning (FSCIT), which facilitates the continual tuning of vision foundation models to learn new classes with only a few samples per class. Unlike traditional FSCIL, FSCIT does not require a large in-distribution base session for initial fully supervised training prior to the incremental few-shot sessions. We conduct extensive evaluations across 16 diverse datasets, demonstrating the effectiveness of CoACT in both FSCIL and FSCIT setups. CoACT outperforms existing methods by up to 5.02% in FSCIL and up to 12.51% in FSCIT for individual datasets, with an average improvement of 2.47%. Furthermore, CoACT exhibits reduced forgetting and enhanced robustness in low-shot experiments. Detailed ablation and sensitivity studies highlight the contribution of each component of CoACT. We make our code publicly available at https://github.com/ShuvenduRoy/CoACT-FSCIL.
@@ 1 Introduction
Large foundation models pre-trained on web-scale unlabeled data exhibit strong generalization capabilities on downstream tasks when fine-tuned with relatively small amounts of labelled data (Radford et al., 2021; Zhou et al., 2022c). However, the immense size of these pre-trained models introduces significant challenges for fine-tuning, especially when working with limited labelled data (Khattak et al., 2023; Yao et al., 2023). Despite recent advancements such as parameter-efficient tuning (Khattak et al., 2023; Gao et al., 2023), it has been observed that fine-tuning the model in a few-shot setting often leads to a decline in the out-of-the-box generalization capability of the foundation model (Roy & Etemad, 2024). Yet, real-world applications not only necessitate learning from a few samples but also demand continual learning of new tasks. Few-shot class-incremental learning (FSCIL) (Masana et al., 2022) is a continual learning setup that extends the scope of few-shot learning by enabling models to adapt incrementally to new tasks while preserving existing knowledge. While few prior works (D’Alessandro et al., 2023; Park et al., 2024) have explored tuning a foundation model in FSCIL setups, existing approaches often struggle with the forgetting of leaned classes and a loss of the foundation model’s generalization capabilities.
@@
Figure 1: Performance comparison on FSCIT with a foundation model.
@@
PriViLege 86.06 75.08 94.10
CoACT (Ours) 84.63 81.19 96.24
Figure 2: Performance comparison to existing methods on tradition FSCIL.
In this work, we propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a novel framework for class-incremental tuning of vision foundation models in few-shot settings. CoACT introduces three novel components: (i) asynchronous contrastive tuning, (ii) controlled fine-tuning, and (iii) consistency-guided incremental tuning. To strike a balance between adaptability to learn new classes and retaining generalizable knowledge of the pre-trained foundation model asynchronous contrastive tuning learns from the first incremental session using a novel asynchronous contrastive loss. Specifically, to provide adaptability to learn new classes, we integrate learnable LoRA modules into the pre-trained encoder and ensure generalization by enforcing consistency between two asynchronous encoders: a student encoder containing the learnable modules and a teacher encoder identical to the pre-trained encoder, updated as the Exponential Moving Average (EMA) of the student. This prevents rapid change in the output distribution of the teacher, which in turn helps reduce overfitting in the student encoder while learning new classes. To further enhance adaptability, we introduce controlled fine-tuning, which is a two-step training protocol for training the first incremental session. First, we train the newly added LoRA modules with a high Learning Rate (LR) for a certain number of epochs and then fine-tune the last few layers of the pre-trained parameters with a relatively lower LR. This also helps balance adaptability with generalizability. Finally, to ensure effective learning of classes in the following incremental sessions while preventing forgetting of previously learned classes and preserving the generalization capabilities of the foundation model at the same time, we introduce a novel regularization technique, consistency-guided incremental tuning. We achieve this by enforcing consistency between the predictions of the learnable encoder in the incremental sessions and the frozen encoder from the first incremental session.
We conduct a comprehensive study on 16 diverse image recognition datasets to investigate the effectiveness of our method in FSCIL. Additionally, we introduce a novel and more challenging setup called few-shot class-incremental tuning (FSCIT), where, unlike FSCIL, we do not assume the availability of a fully-supervised base session. Specifically, in FSCIT, a foundation model is fine-tuned over incremental sessions with a few samples per class for all sessions, including the first session. The datasets include generic objects, fine-grained objects, scenes, satellite images, and texture recognition. Our comprehensive experiments demonstrate that on the FSCIL setup, CoACT outperforms prior methods by up to 5.02%. On our proposed FSCIT setup, CoACT outperforms existing methods by 2.47% on average across 16 datasets, with up to 12.79% performance gain on individual datasets. More importantly, CoACT exhibits reduced forgetting of already learned classes as the number of classes increases. CoACT also exhibit generalization with different sizes of pre-trained encoders and different numbers of shots. We provide detailed ablation studies showing the effectiveness of each component of our method. Overall, our contributions are:
  • We propose CoACT, which consists of three novel components to tune a vision foundation model in a few-shot continual learning setup without losing its generalization capability or forgetting the already learned classes.
  • We introduce a new paradigm of continual learning called Few-shot Class-Incremental Tuning (FSCIT), which aims to tune a foundation model to continuously learn new classes with few samples per class. We also establish several baselines for this new paradigm.
  • Comprehensive experiments show the effectiveness of our method, achieving state-of-the-art (SOTA) on both FSCIT and FSCIL setup. We also show reduced forgetting and effectiveness in very-low-shot settings. Extensive ablation and sensitivity studies show the effectiveness of each of our proposed components.
@@ 2 Related works
2.1 Fine-tuning foundation models.
@@
A number of techniques have recently been proposed to tune foundation models without the need to re-train them from scratch, addressing the growing demand for computationally efficient and parameter-efficient fine-tuning methods. These approaches leverage the pre-trained capabilities of large models, adapting them to downstream tasks without modifying the core pre-trained parameters extensively. Adapter tuning (Houlsby et al., 2019) fine-tunes large pre-trained models to downstream tasks by inserting new learnable layers inside the pre-trained model. On the other hand, prompt-tuning (Lester et al., 2021) and prefix-tuning (Li & Liang, 2021) add learnable prompts with the input embedding for learning the new task without tuning the pre-trained parameters of the model. Low-rank adapters (Hu et al., 2021; Karimi Mahabadi et al., 2021) have been introduced to reduce the computation cost of additional parameters. VPT (Jia et al., 2022) and AdapterFormer (Chen et al., 2022) explored parameter-efficient fine-tuning in the context of vision transformers. A well-known issue of training a foundation model with few labelled samples per class is the overfitting and the loss of the generalization of the foundation model. More recent, CoPrompt Roy & Etemad (2024) proposed a novel regularization technique by enforcing consistency between the pre-trained and learnable encoders to reduce overfitting and improve generalization. Nonetheless, existing methods for tuning foundation models are not designed for continuous tuning of the model since there are no inherent mechanisms to prevent loss of generalization and catastrophic forgetting.
2.2 Few-shot class-incremental learning.
Class-incremental learning is a continual learning process that focuses on continuous learning of new classes while retaining the knowledge of already learned ones (Wang et al., 2023; Smith et al., 2023; Wang et al., 2024; 2022). In practice, machine learning models often need to learn new classes from a few labelled samples per class (Zhou et al., 2022b), while having no access to samples from already learned classes. This scenario has given rise to a new learning task called few-shot class-incremental learning or FSCIL (Tian et al., 2023; Tao et al., 2020). The existing literature on FSCIL can be broadly categorized into two main groups: methods that continuously train both the encoder and classifier over each incremental session  (Cheraghian et al., 2021; Dong et al., 2021; Zhao et al., 2021), and methods that keep the encoder frozen during the incremental learning sessions (Zhu et al., 2021; Shi et al., 2021; Zhang et al., 2021).
As an example of the first group of methods, MgSvF (Zhao et al., 2021) employed a component-wise update strategy to ensure adapting to new classes while preserving knowledge of existing ones. The exemplar relation distillation framework (Dong et al., 2021) constructed and updated a graph of exemplar relationships to facilitate the integration of new classes. SoftNet (Kang et al., 2023) introduced a novel approach for identifying and freezing a sub-network of crucial parameters from the previous session and training the remaining parameters during incremental sessions. While such methods in the first group generally offer greater adaptability, the methods in the second group focus on ensuring stability by maintaining separability in the base classes in the learned embedding space. For example, FACT (Zhou et al., 2022a) introduced virtual prototypes to maximize class embedding separation while preserving their relative positions. SAVC (Song et al., 2023) generated virtual classes during the base session training to maximize separability. Similarly, NC-FSCIL (Yang et al., 2022) pre-assigned optimally spaced prototypes to each base class, promoting diverse and distinct class representations.
Some of the more recent works have explored tuning a foundation in the FSCIL setup. For instance, CPE-CLIP (D’Alessandro et al., 2023) utilized the strong generalization capability of a pre-trained vision-language model to learn new classes with parameter-efficient fine-tuning. SV-T (Qiu et al., 2023) used a semantic-visual guided Transformer to enable few-shot class-incremental learning while reducing overfitting by combining text-based semantic labels with visual labels to improve generalization to new classes. PriViLege (Park et al., 2024) showed that with prompt tuning and knowledge distillation, pre-trained vision language shows strong performances in FSCIL. Nonetheless, these methods operate on the FSCIL setting and still show forgetting issues as they learn new classes.
@@ 3.1 Problem formulation
In FSCIT, a model $\phi(x)$ is continuously trained over $T$ consecutive sessions of new classes. Each session follows an $n$-way, $k$-shot setup, where $n$ represents the number of classes in each session, and $k$ is the number of samples per class. Consequently, the training data for each session $t\in T$ can be represented as $\mathcal{D}_{train}^{t}=\{(x_{i},~{}y_{i})\}_{i=1}^{N_{t}}$, where $x_{i}$ and $y_{i}$ denote the $i^{\text{th}}$ sample and its corresponding label, and $N_{t}$ represents the number of samples in session $t$. We can express $\phi(x)=W^{T}f_{\theta}(x)$, where $f_{\theta}$ is an encoder and $W$ is a linear classifier. In the conventional setup of FSCIL, $f_{\theta}$ is trained on a large labelled in-distribution base session ($\mathcal{D}_{train}^{0}$) in a fully-supervised manner. In FSCIT, we do not assume the availability of such a base session, rendering standard FSCIL techniques incompatible. In this work, we aim to design a method for tuning off-the-shelf vision foundation models to continuously learn new classes in few-shot settings.
Figure 3: Illustration of CoACT. (Left) Training on the first incremental session with asynchronous contrastive tuning and controlled fine-tuning. The student encoder contains learnable LoRA modules, while the teacher is identical to the foundation model but updated as the EMA of the student. Controlled fine-tuning enables the tuning of a subset of the foundation model with reduced LR after certain epochs. (Right) Consistency-guided incremental tuning enforces consistency between the learnable student and the frozen encoder from the first session, providing additional regularization that prevents overfitting and forgetting.
@@ 3.2 Consistency-guided Asynchronous Contrastive Tuning (CoACT)
3.2.1 Asynchronous contrastive tuning.
To strike a balance between adaptability to learn new classes and retaining generalizable knowledge of the foundation model, we introduce asynchronous contrastive tuning as the first component in our framework. This involves fine-tuning the pre-trained model using our novel Asynchronous Contrastive Learning (ACL) approach while incorporating LoRA modules into the model. Let $h_{i}=f_{\theta}^{(i)}(h_{i-1})$ be the output of $i^{\text{th}}$ layer of the pre-trained encoder, $h_{i-1}$ be the output of the $(i-1)^{\text{th}}$ hidden layer of the encoder, and $f_{\theta}^{(i)}$ be the $i^{\text{th}}$ layer of the encoder. With the new learnable LoRA layers, the output of the $i^{\text{th}}$ layer of the network can be represented as $h_{i}^{{}^{\prime}}=f_{\theta}^{(i)}(h_{i-1})+f^{(i)}_{\text{LoRA}}(h_{i-1})$, where, $f^{(i)}_{\text{LoRA}}$ is the $i^{\text{th}}$ LoRA layer added to the pre-trained encoder. For brevity, we denote the encoder with learnable LoRA layers as $f_{\theta^{{}^{\prime}}}$. We can train $f_{\theta^{{}^{\prime}}}$ on $\mathcal{D}_{train}^{1}$ to learn the first session:
@@
However, it has been shown in prior work that cross-entropy alone does not learn a well-separable embedding space (Song et al., 2023) and has a higher over-fitting tendency, especially in a few-shot setting (Roy & Etemad, 2024). To reduce the possibility of overfitting and retain the generalization in the learnable encoder $f_{\theta^{{}^{\prime}}}$, we regularize its output distribution with an asynchronous teacher encoder by maximizing their agreement in the embedding space. Our novelty lies in the asynchronous design of the encoders, where the student encoder contains the learnable LoRA modules, but the teacher encoder is identical to the pre-trained model (without LoRA modules) and learned through the Exponential Moving Average (EMA) of the student $f_{\theta^{{}^{\prime}}}$ as $\theta^{{}^{\prime\prime}}=m\cdot\theta^{{}^{\prime\prime}}+(1-m)\cdot\theta$, where $\theta$ and $\theta^{{}^{\prime\prime}}$ are the parameters of the student (excluding LoRA) and teacher encoders, and $m$ is the momentum parameter. Since the teacher lacks LoRA modules, the EMA update applies only to the student’s base parameters, affecting the teacher layers corresponding to those fine-tuned in the student (Section 3.2.2). Given that the teacher and student encoders differ in their architecture due to the addition of LoRA to the student, learning occurs asynchronously. This asynchronous encoder design and slow-moving update of the teacher through EMA ensures that the predictions from the teacher do not fluctuate. Since the teacher encoder is also initialized from the foundation model, consistency with the teacher effectively regularizes the student from overfitting.
In practice, we maximize the agreement between the embeddings of the student and teacher encoders on all the samples from each class as:
@@
The exclusion of LoRA adapters from the EMA updates prevents rapid fluctuations in the teacher encoder, ensuring it serves as a slow-moving reference that mitigates overfitting and retains the generalization capabilities of the pre-trained foundation model. If LoRA adapters were included, the teacher would change more dynamically, reducing its ability to regularize the student effectively and undermining the balance between adaptability and generalization retention. Maintaining the teacher closer to the pre-trained model prevents the student from drifting too far and acts as a form of regularization that constrains learned representations while preserving prior knowledge.
3.2.2 Controlled fine-tuning.
@@
Controlled fine-tuning is a two-step training protocol to enhance the adaptability of the model by selectively fine-tuning a few layers of the pre-trained encoder. Since the newly added LoRA modules are randomly initialized, we begin by training only the LoRA modules of the student encoder with a higher learning rate for the initial $E_{c}$ epochs of training to prevent the propagation of randomness to the well-trained pre-trained encoder, which could otherwise lead to forgetting of its generalizable knowledge. This is followed by a fine-tuning stage where the last $C_{l}$ layers of the pre-trained encoder are fine-tuned with a reduced LR (scaled by a factor of $C_{f}$). We focus on fine-tuning only the last $C_{l}$ layers, as the later layers of a pre-trained model are responsible for learning domain-specific fine-grained features, whereas the earlier layers are more general and transferable to a wide range of tasks (Neyshabur et al., 2020). Importantly, this fine-tuning is performed under the consistency constraint of asynchronous contrastive learning, ensuring the model learns target-domain features effectively without overfitting or forgetting pre-trained knowledge. This balance between adaptability and generalization enables effective adaptation to new classes while preserving the pre-trained encoder’s foundational capabilities. The two-setup controlled fine-tuning strategy is applied during the first session, while in later sessions, fine-tuning continues with LoRA across all layers, along with fine-tuning the last $C_{l}$ layers.
3.2.3 Consistency-guided incremental tuning.
@@
While the first two modules facilitate tuning the foundation model ($f_{\theta^{{}^{\prime}}}$) during the first session, later sessions also require the retention of previously learned classes. To facilitate this, we propose consistency-guided incremental tuning, which prevents forgetting by regularizing the output distribution of the student $f_{\theta^{{}^{\prime}}}$ when training on the incremental sessions. More specifically, we enforce consistency between the predictions of the student encoder and the frozen student encoder after the first session, effectively discouraging substantial changes in the learned representations of the student. We do not enforce consistency with the frozen encoder from the most recent session (e.g., using the frozen encoder from session $t-1$ for session $t$), since the model may gradually drift away from the pre-trained encoder’s knowledge with cumulative overfitting over the sessions, which can lead to increased forgetting of the foundational generalization capabilities of the model. By anchoring the consistency to the first session’s encoder, CoACT effectively balances the learning of new classes with the retention of pre-trained knowledge, mitigating catastrophic forgetting. In a standard class-incremental learning setup, the encoder updates across multiple sessions as $f_{\theta_{t}}=f_{\theta^{t-1}}+\Delta_{\theta_{t}}$ where $\Delta_{\theta_{t}}$ represents the parameter updates due to learning new classes at session $t$. If is consistency enforced with the encoder from the previous session, $t-1$. As a result, the model incrementally adapts but can also gradually drift away from its initial state, leading to $f_{\theta_{T}}=f_{\theta^{0}}+\sum_{t=1}^{T}\Delta_{\theta_{t}}$. This, in turn, can result in a gradual loss of generalization, where the model drifts away from the foundational encoder. Instead, enforcing consistency with the frozen encoder from the first session $f_{\theta^{0}}$ prevents progressive divergence. This ensures that at each session, $f_{\theta_{t}}=f_{\theta^{0}}+\Delta^{\prime}_{\theta_{t}}$, where $\|\Delta^{\prime}_{\theta_{t}}\|\ll\|\sum_{i=1}^{t}\Delta_{\theta_{i}}\|$.
Let $f_{\theta_{\beta}}$ be the frozen encoder after the first session, and the frozen embedding of this encoder be $p_{i}=f_{\theta_{\beta}}(\mathcal{A}(x_{i}))$. We define our consistency regularizer as: